OpenStudy (anonymous):

I need to determine the probability that x is < or = to 92, when the mean is 100 and the standard deviation is 5

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

convert this to standard normal: \[\frac{x- \mu}{\sigma}\] (92-100)/5=-1.6 -1.6 is the z score. Do you know how to use a standard normal chart?

OpenStudy (anonymous):

No, I don't understand the chart.

OpenStudy (anonymous):

Well there are a few different versions, but the one i typically use shows the probability to the left of each z value. So if you want the probability less than or equal to z=-1.6 you would find z=-1.6 on the chart and read off the value.

OpenStudy (anonymous):

let me find a chart online.

OpenStudy (anonymous):

okay. It was given in multiple choice and the options are .1359, .0215, .1548 and .4452... so if I can just figure out how to get which answer is right, I can have it narrowed down like that.

OpenStudy (anonymous):

ok so then I put in -1.6 on the chart?

OpenStudy (anonymous):

there is a chart. but it only has positive values of z, so you will need to find +1.6 and take 1-answer. 1-.9452=.0548

OpenStudy (anonymous):

did you mean .0548 is a choice?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok do you see how to use that chart I posted?

OpenStudy (anonymous):

yes! thank you so much. I didn't see the chart in my book but I found it in the back.

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

so then a similar problem to that would be the mean of test scores is 500. the standard deviation is 100. what would the cutoff be for the top 10% of scores?

OpenStudy (anonymous):

would I use the chart for that too?